# Large Language Model Can Continue Evolving From Mistakes

### v4

@@ Abstract
As world knowledge evolves and new task paradigms emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings. In practical applications, LLMs often require both continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new task paradigms and acquire necessary knowledge for task-solving. However, it remains challenging to collect CPT data that addresses the knowledge deficiencies in models while maintaining adequate volume, and improving the efficiency of utilizing this data also presents significant difficulties. Inspired by the ’summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient approach for collecting CPT data and continually improving LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To efficiently utilize these CPT data and mitigate forgetting, we design a novel CL training set construction paradigm that integrates parallel CIT and CPT data. Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in accuracy in the best case. Furthermore, additional experiments confirm the potential of combining CEM with catastrophic forgetting mitigation methods, enabling iterative and continual model evolution.
@@ 1 Introduction
With the exponential growth of data and parameter sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and task requirements necessitates ongoing task adaptation and knowledge supplementation for LLMs. Otherwise, LLMs may fall into hallucinations and generate incorrect responses Huang et al. (2023). Therefore, Continual Learning (CL) is proposed to ensure LLMs’ alignment with the evolving world knowledge and demands.
@@
Figure 1: The figure illustrates two potential triggers for poor model performance: (1) Task Paradigm Unfamiliarity, and (2) Lack of task-relevant knowledge. Unfamiliarity with the task paradigm can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations. Instruction tuning addresses only the former.
Wu et al. (2024) categorize Continual Learning into Continual Pre-training (CPT) and Continual Instruction Tuning (CIT). CPT requires a larger volume of training data to enrich LLMs’ knowledge for new domains, while CIT continually fine-tunes LLMs on emerging task-specific instruction data to adapt the models to evolving task paradigms with much smaller training costs and data requirements Razdaibiedina et al. (2023); Gururangan et al. (2020b); Cossu et al. (2022). However, recent work has highlighted the limitations of instruction tuning, suggesting it is merely a superficial alignment process focused on learning interaction styles and LLMs primarily acquire knowledge during the pre-training stage Zhou et al. (2023). This will trigger the bad response shown in Figure 1, where LLMs’ poor performance may result from a lack of requisite knowledge to solve the task rather than unfamiliarity with the task paradigm Ren et al. (2023); Zhang et al. (2023b). CPT is more effective than CIT in addressing this issue.
@@
However, using CPT to supplement the model with the lacking knowledge required for task-solving presents several issues: First, collecting CPT data is challenging, as it is difficult to a priori identify the lacking knowledge while ensuring a sufficiently large data scale. Blind data collection and insufficient data volume may introduce irrelevant information, weaken generalization, and pose overfitting risks Zhang et al. (2024); Gekhman et al. (2024). Second, due to the weak constraints of self-supervised CPT data during parameters updates and the scale gap compared to initial pre-training data, efficiently utilizing CPT data and avoiding catastrophic forgetting remain difficult.
To address these issues, recent work has explored converting CPT data into CIT data, such as transforming it into reading comprehension questions to avoid the problem of insufficient data volume. Cheng et al. (2024) or using highly relevant QA pairs for CIT before CPT to enhance the data utilization efficiency Jiang et al. (2024). However, these methods demand high-quality data and the assistance of oracle LLM, and they overlook the crucial challenge of CPT mentioned above. Inspired by the ‘summarizing mistakes’ learning skill of human students, we argue that a model’s mistakes in ’exams’ reflect its inherent knowledge deficiencies. CPT data that directly addresses these gaps is the most effective for supplementing the model’s knowledge. Currently, the successful utilization of mistakes in constructing CoT data An et al. (2024); Tong et al. (2024) has enhanced the model’s performance in mathematics and reasoning, indicating the potential to leverage mistakes in continual learning.
In this paper, we propose the Continue Evolving from Mistakes(CEM) method, aiming to provide a data-efficient method for collecting CPT data and continually enhancing the performance of LLMs through multiple evaluation and supplementation with mistake-relevant knowledge. Specifically, at the start of each CEM round, we test the model, identify cases of poor performance, and deconstruct them to pinpoint knowledge deficiencies. We then gather abundant supplemental data based on mistake analysis from search engines and online encyclopedias as CPT data to address these gaps. To efficiently utilize these CPT data and mitigate forgetting, we designed a novel CL training set construction paradigm that integrates parallel CIT and CPT data. By integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution.
Our contributions are summarized as follows: (1) We propose the CEM method, which identifies knowledge deficiencies from LLMs’ mistakes and collects abundant relevant CPT data, enabling continual iterative evolution. (2) To assist LLMs in acquiring and comprehending factual knowledge from CPT data and mitigating knowledge forgetting, we propose three strategies for constructing supplemental training set using a parallel paradigm with CIT and CPT data. (3) We conduct experiments on multiple open-source LLMs for two complex knowledge QA tasks. The CEM method consistently demonstrates significant improvements across all scenarios, , achieving up to a 17.00% accuracy enhancement. (4) We verify that CEM can be combined with methods to mitigate catastrophic forgetting, such as replay-based methods Chaudhry et al. (2019); Wu et al. (2021), ensuring suppression of catastrophic forgetting over multiple iterations.
@@ 2 Related Work
Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks. A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al. (2023). However, studies have shown this process causes catastrophic forgetting of the learned knowledge and problem-solving skills from previous tasks Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3.2 CEM Method
Step 3: Obtaining Supplemental Corpus from Multiple Data Sources. Based on the implicit information of knowledge deficiencies contained within questions incorrectly answered by the LLM, we collect supplemental corpus from multiple data sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are characterized by high relevance matching and diverse content. We search for questions the LLM answers incorrectly. The search engine ranks results by relevance, and we select the top-k results to access their detailed pages, extracting the main textual content as supplemental corpus. To prevent the LLM from merely memorizing answers, we filter out results that contain both the original question and its answer. Additionally, we optimize the corpus crawling strategy to exclude irrelevant information and spam, such as advertisements.
In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 6 in Appendix shows the process of corpus collection from Bing.
(2) Online Encyclopedia: Compared to search engines, online encyclopedias provide more credible and professional results, despite lower hit rates. Since encyclopedias are typically organized by keyword entries, we conduct entity extraction from the questions and options, using the extracted entity keywords for retrieval. The main text of the retrieved results serves as the supplemental corpus.
In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction. Figure 7 in Appendix shows the process of corpus collection from Wikipedia.
@@ 3.3 Supplemental Training Data
(1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data.
@@
(2) Normative Corpus($C_{\underline{N}}$): To enable the LLM to answer multiple-choice questions in a normative format, we sample questions independent of the test set and process them into the format shown in Table 1 to build the Normative Corpus. Prior to supplemental training, we finetune the LLM using the Normative Corpus (i.e., pre-finetuning), which facilitates extracting answers from the LLM’s responses to calculate initial performance and allows for the collection of incorrectly answered questions. To maintain this ability, we incorporate a small proportion of the Normative Corpus during supplemental training.
@@
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions independent of the test set and the Normative Corpus. These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
@@
The responses of the LLM tend to be diverse under high temperature settings Caccia et al. (2018). During repeated QA testing, the LLM may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
@@
(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered as the General Corpus to preserve the model’s general capabilities on other tasks.
@@ 4.1 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024).
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities. Appendix Table 6 shows the details of the data division and usage.
@@ 4.2 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion. Our instruction fine-tuning base models include CuteGPT-13B [^2] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al. (2023), ChatGLM-6B [^3] https://huggingface.co/THUDM/chatglm3-6b Du et al. (2022), and Qwen-7B-Chat [^4] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al. (2023). These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
@@ 4.3 Metrics
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{i}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
@@ 5 Experiments
We construct supplemental training sets and control groups for the experiments, following the strategies described in Section 3.3. To distinguish between models trained on different supplemental training sets, we use the following naming conventions:
INIT represents the model that has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
CEM-P: P denotes Plain, representing the model trained with the supplemental training set constructed by the Plain Strategy.
@@
We append suffixes to indicate the data sources of the Supplemental Corpus: ‘Wiki’, ‘Bing’ and ‘Mix’. ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated. For example, ‘E-Wiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
@@ 5.1 Main Results
We used the Plain Strategy to construct supplementary training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM. The main results are shown in Figure 3. Our observations are summarized as follows:
@@
Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two tasks. This indicates that the CEM method effectively addresses the model’s knowledge deficiencies, even with the small training set and readily available supplemental data. For instance, the ChatGLM-6B improves by up to 16.08% on the CMMLU task.
@@
Firstly, Extractive and Review Corpus significantly enhance LLM performance. CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics. Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric. The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source. Results in Figure 4 show no significant difference in CEM effectiveness whether using Wikipedia, Bing, or a combination of both. Indeed, R-Mix has lower accuracy than R-Wiki and R-Bing despite having more data. This may be due to redundancy from overlapping sources. Additionally, an excess of supplemental corpus can dilute the proportion of CIT data, reducing their effectiveness in mitigating catastrophic forgetting and enhancing knowledge extraction, and increasing the risk of overfitting.
Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of Review Corpus. Table 5 in Appendix shows that R0, R0.5, and R1 have similar effects on LLM performance. This suggests that reviewing the knowledge in any format consolidates it. Therefore, unless specified, R will refer to the R0.5 in the following.

### v5

@@ Abstract
As world knowledge evolves and new task schemas emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings. LLMs typically require continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new tasks and acquire essential knowledge. However, collecting sufficient CPT data while addressing knowledge gaps remains challenging, as does optimizing the efficiency of utilizing this data. Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To enhance data utilization and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT data. Experiments demonstrate that CEM significantly enhances model performance and continual evolution. The code and dataset are available in the GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
@@ 1 Introduction
With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs. Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al. (2023). Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial. Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al. (2024).
@@
Figure 1: Two potential triggers for poor model performance: (1) Task Schema Unfamiliarity, and (2) Lack of Task-relevant Knowledge. Unfamiliarity with the task schema can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations. Instruction tuning has been shown to be effective for addressing the former, but poor for the latter Gekhman et al. (2024a); Zhou et al. (2023).
Despite CIT’s smaller data size requirement and lower risk of overfitting compared to CPT, it is limited in new knowledge injection, as Zhou et al. (2023) highlight that instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and does not effectively help the model acquire new knowledge. Thus, it cannot address all the issues shown in Figure 1, where LLMs’ poor performance may stem from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al. (2023); Zhang et al. (2023b). In contrast, CPT is more effective in addressing the lack of knowledge.
@@
However, using CPT to supplement the model with lacking knowledge still presents several challenges. First, collecting CPT data is challenging as identifying the lacking knowledge a priori while ensuring a sufficiently large data size is difficult. Blind data collection and insufficient data volume may introduce irrelevant or redundant information, weaken generalization, and increase overfitting risks Zhang et al. (2024); Gekhman et al. (2024b). Second, efficiently utilizing CPT data and mitigating degradation are difficult due to weak constraints on parameter updates, leading to larger parameter shifts and a data size gap compared to the initial pre-training.
To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al. (2022); Jin et al. (2022). However, these methods lack focus and efficiency and fail to dynamically update model deficiencies. Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al. (2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al. (2024) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data. Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation. To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’. We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies. Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes for directly address these deficiencies. To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data. In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution. The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 17.00% accuracy enhancement.
@@ 2 Related Work
Task-incremental Continual Instruction Tuning (CIT) aims to continually fine-tune LLMs on new task-specific instructions. While a straightforward approach involves generating new instruction-tuning data and fine-tuning LLMs directly Wang et al. (2023), this method can lead to catastrophic forgetting of previously learned knowledge and skills Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3.2 CEM Method
Step 3: Obtaining Supplemental Corpus from Multiple Data Sources. We first employ the lightweight LLM, Qwen2-7B-instruct [^2] https://huggingface.co/Qwen/Qwen2-7B-Instruct Yang et al. (2024), to identify knowledge points involved in questions erroneously answered by the tested LLM. For further details, see Appendix A. The extracted knowledge points are then utilized as keywords to aggregate supplemental corpora from diverse sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are known for high relevance matching and diverse content. We search for extracted knowledge points using Bing and select the top-k results based on relevance, scraping the main textual content as a supplemental corpus. To prevent LLMs from simply memorizing answers, we filter out results that contain both the original question and its answer. Moreover, we use newspaper3k [^3] https://github.com/codelucas/newspaper and readability [^4] https://github.com/buriy/python-readability to exclude irrelevant and spam content.
(2) Online Encyclopedia: Compared to search engines, online encyclopedias offer more credible and professional results, despite having lower hit rates. Since encyclopedias are typically organized by keyword entries, we employ the state-of-the-art massive text embedding model, Conan-embedding-v1555https://huggingface.co/TencentBAC/Conan-embedding-v1 Li et al. (2024b), to generate vector embeddings for Wikipedia entry titles. FAISS [^6] https://github.com/facebookresearch/faiss Douze et al. (2024) is then used to build an efficient vector index, enabling searches based on inner product similarity. By calculating the similarity between input keyword embeddings and the FAISS-indexed entry embeddings, we retrieve the top-k results that exceed a predefined similarity threshold, serving as the supplemental corpus.
Appendix C presents the experimental parameters and processes.
@@ 3.3 Supplemental Training Data
(1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data. We employ Qwen2-72B-Instruct [^7] https://huggingface.co/Qwen/Qwen2-72B-Instruct Yang et al. (2024) for entailment checking to verify whether the Supplemental Corpus contains information that helps the model answer correctly. Details are provided in the Appendix B.
@@
(2) Normative Corpus($C_{\underline{N}}$):
To enhance the LLM’s understanding of task schemas and formats, task-relevant instruction fine-tuning is commonly applied. For QA tasks, we sample questions from the training set and format them as shown in Table 3 to build the Normative Corpus, enabling the LLM to answer multiple-choice questions in a standardized manner. Before supplemental training, we fine-tune the LLM using the Normative Corpus (i.e., pre-finetuning). A small portion of the Normative Corpus is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
@@
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions from the training set and format them as shown in Table 4 to build the Extractive Corpus. We employ Qwen2-72B-Instruct for entailment checking to filter the corpus. A minimal amount of Extractive Corpus data yields effective results, ensuring manageable inference costs. Details are provided in the Appendix E.
@@
The responses of the LLM tend to be diverse under high temperature settings Caccia et al. (2018). During repeated QA testing, the LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
@@
(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^8] https://huggingface.co/datasets/Aeala/ShareGPT_
Vicuna_unfiltered
as the General Corpus to preserve the model’s general capabilities on other tasks and instruction-following ability.
@@ 4 Experimental Setup
4.1 Training Setup
We train the models on 2 A800 GPUs using ZeRO (Rajbhandari et al., 2020) stage 1 from DeepSpeed (Rasley et al., 2020). We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 16 with a maximum sequence length of 2,048 and 1 training epoch.
We explore the impact of full fine-tuning, LoRA fine-tuning Hu et al. (2022), and single-stage versus multi-stage training methods. For details, see the Appendix F.
@@ 4.2 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a).
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities. Appendix G shows the details of the data division and usage.
@@ 4.3 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs include CuteGPT-13B [^9] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al. (2023), ChatGLM-6B [^10] https://huggingface.co/THUDM/chatglm3-6b Du et al. (2022), and Qwen-7B-Chat [^11] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al. (2023). These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
@@ 4.4 Metrics
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{k}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
@@ 5 Experiments
We construct supplemental training sets and control groups for experiments as outlined in Section 3.3. To differentiate models trained on various supplemental sets, we employ the following naming conventions:
INIT represents the model pre-finetuned solely with the Normative Corpus.
CEM-P: ‘P’ for Plain, indicates the model trained using the supplemental training set constructed by the Plain Strategy.
@@
Suffixes ‘Wiki’, ‘Bing’, and ‘Mix’ indicate the data sources of the Supplemental Corpus, with ‘Mix’ signifying a combined set from Bing and Wikipedia. For instance, ‘EWiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
@@ 5.1 Main Results
We use the Plain Strategy to construct supplemental training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM. The main results are shown in Figure 3. Our observations are summarized as follows:
@@
Firstly, after supplemental training, a marked improvement in accuracy is noted across all models for both tasks. This suggests that the CEM method effectively addresses the model’s knowledge deficiencies, despite the limited training set and readily accessible supplemental data. For example, the ChatGLM-6B exhibits an increase of up to 16.08% in accuracy on the CMMLU task.
@@
Firstly, Extractive and Review Corpus significantly enhance LLM performance. CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics. Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric.
The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
Secondly, a Supplemental Corpus from multiple sources does not invariably enhance the CEM method more than a single source. Results in Figure 4 indicate no substantial difference in CEM effectiveness, whether using Wikipedia, Bing, or their combination. Notably, RMix records lower accuracy compared to RWiki and RBing, despite a larger dataset. This may arise from data redundancies. Moreover, an excessive supplemental corpus can reduce the proportion of CIT data, undermining its role in mitigating catastrophic forgetting, improving knowledge extraction, and increasing the risk of overfitting.
Thirdly, proportion factor $\boldsymbol{\alpha}$ has limited impact on the effectiveness of Review Corpus. Table 9 shows that R0, R0.5, and R1 have similar effects on LLM performance. This suggests that reviewing the knowledge in any format consolidates it. Therefore, unless specified, R will refer to the R0.5 in the following.